Purchase Price is $24,000
Interest rate of their 36 month loan is 2.9%
Each delivery will generate $20 is revenues, but will cost $16 in wages, gas and vehicle wear
They anticipate completing 50 deliveries each day
Assuming 30 days work each month; what is the anticipated monthly profit for this truck during the loan
repayment period. What are the profits for the remaining 4 years of anticipated remaining useful life of the truck?
Excess cash can be saved in a standard interest-bearing account earning 0.5%
Their risk adjusted discount rate is 6%
They pay taxes on their net profits (after all expenses) at a 25% rate.
Consolidate the monthly net profits into annual cash flows. What are the:
PBP
DPBP
NPV
IRR
MIRR
PI
Answer #1

Profit per delivery = 20-16 = $4

Profit per day = $4 x 50= $200

Profit per month = $200 x 30 = $6000

Therefore annual cash flow = $6000 - 0.029 x$ 24,000 = $5,304

Annual net cash flow after tax = $5,304 - 0.25 x $5,304 = $3,978

1. Payback period = 24,000/3,978 = 6.03 years

2. Profit for the remaining four years of life after loan repayment will be $6,000 - 0.25 x 6,000 = $4,500

Discounted value of cash flows = $3,978 x PVFA 6%, 3 years, + $4,500 x PVFA 6%, 4 years

= $10,633 + $13,091 = $23,724

3. NPV is negative at $260

4. IRR and MIRR can be calculated in excel

5. PI is 23,724/24,000 = 0.9885